### v2

# sPhinX: Sample Efficient Multilingual Instruction Fine-Tuning Through N-shot Guided Prompting

Abstract
Despite the remarkable success of LLMs in English, there is a significant gap in performance in non-English languages. In order to address this, we introduce a novel recipe for creating a multilingual synthetic instruction tuning dataset, sPhinX, which is created by selectively translating instruction response pairs from English into 50 languages. We test the effectiveness of sPhinX by using it to fine-tune two state-of-the-art models, Phi-3-small and Mistral-7B and then evaluating them across a comprehensive suite of multilingual benchmarks that test reasoning, question answering, and reading comprehension. Our results show that Phi-3-small and Mistral-7B fine-tuned with sPhinX perform better on an average by 4.2%pt and 5%pt respectively as compared to the baselines. We also devise a strategy to incorporate N-shot examples in each fine-tuning sample which further boosts the performance of these models by 3%pt and 10%pt respectively. sPhinX outperforms other multilingual instruction tuning datasets on the same benchmarks along with being sample efficient and diverse, thereby reducing dataset creation costs. Additionally, instruction tuning with sPhinX does not lead to regression on most standard LLM benchmarks.
1 Introduction
Large Language Models (LLMs) have been shown to perform several tasks exceptionally well in English, however, performance in some non-English languages still lags behind Ahuja et al. (2023); Asai et al. (2023). Further, the gap between the performance of Large Language Models (LLMs) and Small Language Models (SLMs) is more pronounced Ahuja et al. (2024) on
non-English languages. Cui et al. (2023) and Balachandran (2023) follow the approach of training the models further on datasets for specific languages, however this can lead to catastrophic forgetting and hurt the performance on English Zhao et al. (2024). Several techniques have been proposed to bridge this gap, such as incorporating better pre-training data in various languages and enhancing base tokenizers Xu et al. (2024); Dagan et al. (2024). However, most of these changes need to be implemented at the pre-training stage which demands extensive data and computational resources, making it practically unfeasible in many scenarios Brown et al. (2020). Consequently, the most well-studied technique involves fine-tuning models for specific languages and tasks. Instruction fine-tuning (IFT) has become a popular technique for enhancing the performance of language models in specific languages. This method combines the benefits of both the pretrain–fine-tune and prompting paradigms Wei et al. (2021).
Sample diversity is crucial for instruction tuning in multilingual datasets. Many recent datasets are created by translating English content into other languages or using self-instruct techniques from seed prompts Li et al. (2023); Taori et al. (2023), both of which can limit diversity. Machine translation can lead to loss of semantic information Baroni and Bernardini (2006), while self-instruct methods may produce repetitive and homogeneous samples Wang et al. (2022). This underscores the importance of having a dataset with diverse set of samples.
Aggarwal et al. (2024) evaluate several models fine-tuned using Parameter Efficient fine-tuning and find that there is a gain in multilingual performance across several SLMs for many low-resource languages, with some high-resource languages performing worse after fine-tuning. However, the performance gains often do not match the performance of larger models, such as GPT-4 and Gemini and can be inconsistent across languages. Hence, there is a need to study instruction tuning for better multilingual performance in SLMs.
In this paper, we present a novel recipe for creating a multilingual synthetic instruction tuning dataset, sPhinX.  It comprises 1.8M instruction-response pairs in 51 languages, derived by selectively translating the Orca instruction tuning dataset Mukherjee et al. (2023) using GPT-4 Achiam et al. (2023). We assess the effectiveness of sPhinX by fine-tuning two models — Phi-3-small and Mistral-7B — across a range of evaluation benchmarks that test various language model capabilities like reasoning, question answering, and reading comprehension. We compare models fine-tuned on sPhinX with those using other synthetic multilingual instruction tuning datasets like Aya Üstün et al. (2024), Multilingual Alpaca Taori et al. (2023), and Bactrian Li et al. (2023) and observe significant performance gains.
The contributions of this paper are as follows:
  • We introduce a novel approach to generate synthetic data for multilingual instruction tuning by selectively translating the Orca dataset with the assistance of GPT-4 (§3.1)
  • We devise LAnguage-Specific N-shot Guided Instruction fine-tuning (LANG) strategy for enhancing the multilingual capabilities of LLMs (§4.2)
  • We conduct an extensive number of experiments on different multilingual instruction tuning datasets to test generalizability in multilingual settings (§5)
  • We also conduct an in-depth analysis where we find sPhinX to be more sample efficient and diverse across languages (§3.3)
2 Related Work
2.1 Multilingual LLMs
Recently, there has been there has been interest in developing and improving SLMs such as LLaMA Touvron et al. (2023), Mistral Jiang et al. (2023), Phi3 Abdin et al. (2024), and Gemma Team et al. (2024). In a relatively brief period, researchers have developed numerous monolingual and multilingual versions of these foundational models, either by pre-training from scratch on specific languages or groups of languages, or by further fine-tuning on them Nguyen et al. (2023); Gala et al. (2024); Balachandran (2023); Cui et al. (2023); Zhang et al. (2023); Qin et al. (2024). Prior work has shown that smaller models show promising results in high-resource languages but perform worse on low-resource ones Ahuja et al. (2024) compared to larger models. To address this, researchers have proposed using multilingual corpora for pre-training, expanding the vocabulary Wang et al. (2019) and continual training Liu et al. (2023), though this increases training costs significantly. fine-tuning for specific tasks, such as translation, has been explored but often struggles with generalization  Mishra et al. (2021). In contrast, instruction tuning has shown to improve generalization to unseen tasks, enhancing the models’ ability to understand and respond to human instructions  Mishra et al. (2021); Ouyang et al. (2022). This makes the models more versatile and capable of handling a wide variety of tasks.
2.2 Multilingual Instruction fine-tuning
Early studies focused on fine-tuning pre-trained models on a variety of languages through data augmentation for a single task Hu et al. (2020); Longpre et al. (2021); Asai et al. (2022). Currently, the approach has shifted on fine-tuning these models using a wide-variety of tasks Longpre et al. (2023); Ouyang et al. (2022). Models such as BLOOMZ Muennighoff et al. (2022) and mT0 Muennighoff et al. (2022) makes significant strides in improving the multilingual performance of decoder-based models Ahuja et al. (2023). There have been multiple multilingual instruction datasets and models proposed such as Bactrian Li et al. (2023), AYA Üstün et al. (2024), polyLM Wei et al. (2023b) after BLOOMZ and mT0. However, these models still do not perform as well as English on other languages, with the gap being particularly large for low-resource languages and languages written in scripts other than the Latin script  Ruder et al. (2021); Ahuja et al. (2023); Asai et al. (2023); Ahuja et al. (2024). In this work, we aim to narrow the performance gap by introducing a strategy for creating datasets for multilingual instruction tuning and recipes for fine-tuning, which we will discuss in the following sections.
Figure 1: The above figure describes translating using a Translation API vs Selective Translation
2.3 Multilingual Synthetic Data Generation
Most instruction-tuning datasets across multiple languages typically focus on general tasks rather than specific reasoning capabilities. While datasets like Orca Mukherjee et al. (2023) and Orca 2 Mitra et al. (2023) exist in English, they highlight a prevalent issue: current methods often prioritize style imitation over leveraging the reasoning abilities found in large foundation models (LFMs). The Orca dataset addresses this by imitating rich signals from GPT-4, including explanation traces and step-by-step thought processes Wei et al. (2023a), guided by assistance from ChatGPT. In order to create multilingual datasets, researchers commonly use translation APIs or LLMs to translate English-specific datasets into target languages. For example, the Bactrian dataset Li et al. (2023) translates Alpaca and Dolly instructions into 52 languages using the Google Translator API and generates outputs with GPT-3.5 turbo. However, these translated datasets often struggle to encode semantic information effectively Baroni and Bernardini (2006). Our dataset approach aims to tackle these challenges by selectively translating only the essential portions of multilingual inputs. This strategy not only preserves semantic information but also accommodates diverse linguistic contexts, thereby enhancing the overall quality and applicability of instruction-tuning datasets across languages.
3 sPhinX Dataset
In this section, we describe our dataset construction methodology (§3.1), dataset filtering and cleaning pipelines (§3.2), and exploratory analysis (§3.3) for determining diversity of our data in comparison to other instruction tuning datasets.
Figure 2: Some examples of input queries and its counterpart existing in the hindi version of the Multialpaca dataset and if it was generated using the Selective Translation strategy
3.1 Dataset Construction
Mukherjee et al. (2023) illustrate how smaller models can replicate the reasoning processes of LLMs and learn from detailed signals like as explanation traces, step-by-step thought processes, and other intricate instructions obtained from data annotated by ChatGPT and GPT-4111GPT-4 inference hyper-parameters in Azure OpenAI interface set as: temperature=0.0, top_p=1, frequency_penalty=0, presence_penalty=0, stop=[”¡—im_end—¿”, ”¡—im_start|¿”]. Their dataset consists of 5M ChatGPT and 1M GPT-4 generated instruction-response pairs. Inspired by this work, we utilized 1M GPT-4 generated instruction-response pairs of their dataset and construct our dataset along similar lines by selectively translating them into 50 different languages with the help of GPT-4. We categorize them into three groups: high-resource, mid-resource, and low-resource languages as outlined in Table 11. For high resource languages, we randomly sample 100k instruction-response pairs from the Orca 1M dataset and generate the responses from GPT-4 using Selective Translation as shown in Figure 1. Similarly, we leverage the same strategy for medium and low resource languages by sampling 50k and 25k pairs respectively. Although GPT-4 performs competitively with commercial translation systems (Google Translate & Bing Translate) it still lags behind on medium and low resource languages Jiao et al. (2023); Hendy et al. (2023). Furthermore, as highlighted in Chang et al. (2023), fine-tuning with a large set of samples from medium and low-resource languages might lead to catastrophic forgetting of high-resource languages. Therefore, we deliberately create fewer samples for medium and low-resource languages than for high-resource ones.
A fundamental problem with using an off-the-shelf translation API is the lack of semantic and task awareness, in addition to translationese Baroni and Bernardini (2006)), which can result in poor quality training data. Consider for example the task of Machine Translation as part of the instruction, wherein the language of the source sentence should be retained. However, an off-the-shelf API, without task awareness, would translate it, resulting in an ill defined instruction. To mitigate this issue, we prompt GPT-4 to selectively translate the instructions, so that the tasks are translated into the appropriate language without changing the semantic meaning. Figure 2 illustrates this with concrete examples. The first example demonstrates the aforementioned translation inconsistency issue for an instruction asking for a French equivalent of an English phrase. The second example demonstrates the consequence of direct translations in the MultiAlpaca dataset: wherein the translation of the task input results in the task being ill-defined based on the instructions. As demonstrated, our proposed Selective Translation method is able to keep the semantic information of the task intact while translating the instructions. For the exact prompt, please refer to Figure 6 in the Appendix.
3.2 Dataset Filtering and Cleaning
Input: listOfSentences: list of strings
1 Function dataFilter(listOfSentences):
2       englishWords $\leftarrow$ set of English words from NLTK;
3       foreach sentence in listOfSentences do
4             cleanedSentence $\leftarrow$ replace all punctuations, digits, and single characters with a single space;
5             cleanedSentence $\leftarrow$ replace all sequences of whitespace with a single space;
6             wordCount $\leftarrow$ 0;
7             foreach word in cleanedSentence do
8                   if word.lower() in englishWords then
9                         wordCount $\leftarrow$ wordCount + 1;
10
11
12
13            content $\leftarrow$ wordCount / len(cleanedSentence);
14             if content $>$ 0.90 then
15                   remove sentence from listOfSentences
16
17
Algorithm 1 Data Filtering Algorithm
Figure 3: t-SNE Visualisation of 1000 samples equally distributed in 10 languages
After generating these translations, we analyze how many instruction-assistant pairs were unsatisfactorily translated, such as incomplete or missing translations, by identifying the English content within the instances. We use the NLTK [^2] https://www.nltk.org/ library to identify the fraction of English words in the sample. The NLTK corpus contains approximately 240,000 English words, making it suitable for our use case. We manually analyzed various proportions of English content in samples of each language and found that the samples were satisfactory when the English content was within 90% for all languages. If the English content exceeded 90% in a language that uses a different script, those samples were removed from our dataset as shown in our data filtering algorithm here. Finally, we manually inspected a small sample of the dataset to ensure the quality of the translations, and found them to be of good quality.
After filtering, the final dataset comprised of 1.8 million samples in 51 languages, divided into three subsets: Train, Test, and Few-shot. We partitioned each language’s dataset to ensure the Test and Few-shot sets contained 2000 and 1000 samples respectively, with the Train set consisting the remaining data. This approach guarantees consistent distributions across languages in the Test and Few-shot sets, ensuring equitable representation regardless of training distribution. The train, test, and few-shot sets are in the ratio of 92 : 5.3 : 2.7.
3.3 Data Analysis
A primary issue that we hypothesize (and find preliminary evidence for from basic analysis of prior Multilingual IFT datasets) is the lack of diversity of samples. In order to validate this issue, we use t-Distributed Stochastic Neighbor Embedding (t-SNE) Cai and Ma (2022) to visualize the structure of both prior multilingual IFT datasets as well as sPhinX. We use the model paraphrase-multilingual-MiniLM-L12-v2 from huggingface’s sentence-transformers Reimers and Gurevych (2020), Reimers and Gurevych (2019) library [^3] https://huggingface.co/sentence-transformers/paraphrase-multilingual-MiniLM-L12-v2 to generate the embeddings. The hyperparameters were set as follows: learning rate at 200, perplexity at 30, and number of iterations (n_iter) at 1000. Concretely, we use 1000 samples equally distributed in 10 languages. Figure 3 demonstrates that sPhinX demonstrates greater diversity compared to prior datasets. Particularly, Bactrian and Multilingual Alpaca appear less diverse, which we hypothesize might be due to their generation method involving translations of the same set of task limited English samples into different languages. By virtue of using a more diverse seed set (Mukherjee et al., 2023), we circumvent this issue by design. Furthermore, one notable differentiating factor of sPhinX is the substantial amount of code-switched data the inherent nature of Selective Translation. This in turn enhances the diversity of samples across multiple languages. Finally, we also observe the Aya dataset to be more diverse than other prior counterparts, though it still shows less diversity across task instructions.
Dataset Average Token
Length/Sample
Aya 2240
Bactrian 2465
Multilingual Alpaca 1620
sPhinX-0s 544
sPhinX 3100
Table 1: Average Token Length in each dataset
4 Experiments
4.1 Setup
Base Models: We use Mistral-7B  [^4] We specifically use the v1.0 base model from here. Jiang et al. (2023) and Phi-3-small Abdin et al. (2024) base model variants and fine-tune them on different datasets based to improve multilingual performance.
Datasets: Apart from the sPhinX dataset, we use Bactrian Li et al. (2023), MultiAlpaca Wei et al. (2023b) and Aya Singh et al. (2024) instruction datasets for comparative evaluation §3.1.
  • Bactrian Li et al. (2023) is a machine translated dataset of the original alpaca-52k Taori et al. (2023) and dolly-15k Conover et al. (2023) datasets into 52 languages. The instructions for this dataset were translated using a Translation API and then GPT-3.5-Turbo was prompted to generate outputs. We fine-tune our models on the complete dataset consisting of 3.4M instances.
  • MultiAlpaca Wei et al. (2023b) is a self instruct dataset which follows the same approach as the English-only Alpaca by translating the seed instructions to 11 languages and then using GPT-3.5.-Turbo to generate responses. We fine-tune our models on the complete set of the dataset consisting of 500k datapoints.
  • Aya Singh et al. (2024) consists of human-curated prompt-completion pairs in 65 languages called the Aya Dataset. It also consists of an aggregation of 44 monolingual and multilingual templated instruction datasets and 19 translated datasets ranging over 114 languages. The total size of the Aya dataset is around 513M instances. Due to the in-feasibility of fine-tuning our models on the entire dataset, we sampled it down to 2.7M instances, ensuring parity with our sPhinX dataset by selecting equal numbers of samples for each language in our subset.
Evaluation: We evaluate [^5] Evaluation prompts and other details in Appendix §A.1 our fine-tuned models along with the available Instruction fine-tuned model variants of Mistral-7B and Phi-3-small (IFT [^6] We take the Mistral-7B instruction-tuned variant from here and Phi-3-small variant from here.) on XCOPA Ponti et al. (2020), XStoryCloze Lin et al. (2022), XWinograd Muennighoff et al. (2023)Tikhonov and Ryabinin (2021), Belebele Bandarkar et al. (2023), and XQuAD Artetxe et al. (2020) using the language model evaluation harness Gao et al. (2023).
  • XCOPA: A causal commonsense reasoning dataset in 11 languages, evaluated in a 4-shot prompt setting.
  • XStoryCloze: A professionally translated version of the English StoryCloze dataset Mostafazadeh et al. (2017) in 10 languages, evaluated in a 4-shot prompt setting.
  • Belebele: A parallel reading comprehension dataset across 122 languages, with evaluation on a subset of 14 languages in a 0-shot prompt setting.
  • XQuAD: A QA dataset consisting of professional translations of a subset of SQuAD into 10 languages, evaluated in a 3-shot prompt setting due to context window limitations.
  • XWinograd: A collection of Winograd Schemas in six languages for cross-lingual commonsense reasoning, evaluated in a 0-shot setting.
4.2 Fine-Tuning Methodology
Inspired from  Longpre et al. (2023)’s strategies to instruction tune a model, we devise Language-Specific N-shot Guided Instruction fine-tuning (LANG). With this approach, we augment a training example by prepending $N$ number of samples of same language as that of the original training example randomly selected from the corresponding few shot set. This augmented training example is used for instruction tuning our models. Suppose a training example in language $l$ is a pair of $Instruction({I}_{\text{train}}^{l})$ and $Response({R}_{\text{train}}^{l})$. We need to prepend $N$ number of shots represented by $({I}_{\text{fewshot}_{1}}^{l},{R}_{\text{fewshot}_{1}}^{l})$, $({I}_{\text{fewshot}_{2}}^{l},{R}_{\text{fewshot}_{2}}^{l})$, …, $({I}_{\text{fewshot}_{N}}^{l},{R}_{\text{fewshot}_{N}}^{l})$. For every training example, $N$ is chosen with a probability $p$ as defined in Table 3. $\mathcal{T}$ is the instruction tuning templating function which takes $Instruction$ and $Response$ and transforms it to user-assistant format by adding the special tokens. Thus, the final training example is:
$(\bigoplus_{i=1}^{N}\mathcal{T}(I_{\text{fewshot}_{i}}^{l},R_{\text{fewshot}_{i}}^{l})\oplus\mathcal{T}(I_{\text{train}}^{l},R_{\text{train}}^{l}))$
where $N$ is chosen based on the probability distribution $P(N)$. $\bigoplus$ represents combining all few-shot examples, while $\oplus$ indicates concatenating the main training example with the aggregated few-shot examples. Once $N$ is selected, an equivalent number of few-shot examples is sampled uniformly at random from the few shot set. The maximum context token length for training the models is set at 8192. To ensure augmented samples fall within this range, we assign more weights to $N$=0 and $N$=1. This strategy is consistently applied when instruction tuning other datasets for both Phi-3-small and Mistral-7B base model. The average token length of samples in each datasets as per Phi-3-small model tokenizer is shown in Table 1. To assess the effectiveness of each instruction tuning dataset on an equal scale, we conducted a comparative analysis of model performance, fine-tuning each on approximately 8 billion tokens from each dataset using LANG strategy.
4.3 Hyperparameters and Training Setup
We used 5 nodes with each node containing 8 A100 GPUs with 80GB VRAM. These nodes communicated with each other using InfiniBand  [^7] https://network.nvidia.com/pdf/whitepapers/IB_Intro_WP_190.pdf. We use DeepSpeed Rasley et al. (2020) to do distributed fine-tuning over these GPUs. We use the same hyperparameters (Table 2) to fine-tune both Mistral-7B and Phi-3-small models.
5 Results
Figure 4: Performance of Mistral-7B and Phi-3-smallwhen instruction-tuned on 8B tokens across various datasets on different benchmarks.
Figure 5: Overall performance for both Mistral-7B and Phi-3-small on our dataset and fine-tuning recipe. Legend: Base_Model: Base Model, IFT: Instruction Fine-Tuned variant of the Base Model available publicly, Bactrian: Base model trained on Bactrian dataset, mAlpaca: Base Model trained on mAlpaca dataset, Aya: Base Model trained on a subset of the Aya dataset, Sphinx-0s: Base Model trained in a 0-shot fashion on the sPhinX dataset, Sphinx: Base model trained using LANG on sPhinX dataset
We evaluate the reasoning, question answering, and reading comprehension abilities of the Phi-3-small and Mistral-7B models, instruction-tuned on different multilingual datasets, using various benchmarks and find that fine-tuning on sPhinX provides an improvement of 4.2%pt [^8] A percentage point (pt) is the unit for expressing the absolute difference between two percentage values and 5%pt respectively ( language-wise results are in the Appendix §A.2). Additionally as observed in Figure 4, the sPhinX dataset significantly enhances the multilingual performance of the Phi-3-small and Mistral-7B model compared to other datasets even when finetuned on equal number of tokens. Furthermore, during instruction tuning on 8B tokens, the models encountered fewer training samples for sPhinX due to its higher average token length per sample, as illustrated in Table 1.
Due to the code-mixed nature of the instruction along with CoT reasoning explanations, a single sample of sPhinX is notably richer as compared to its counterparts from the other datasets. Consequently, even with fewer samples (keeping the number of the tokens same), models trained on sPhinX achieve better performance; thereby demonstrating the per sample efficiency of sPhinX.
6 Ablation: Improvements from LANG
To demonstrate the effectiveness of our LANG strategy, we also instruction-tuned the models on sPhinX without any pre-pended shots, referring to this as sPhinX-0s. As shown in Figure 5 (with detailed results in Appendix §A.2), models fine-tuned on sPhinX especially Mistral-7B exhibits superior performance compared to its counterparts fine-tuned on other datasets across all benchmarks. Moreover, fine-tuning both Phi-3-small and the Mistral-7B on sPhinX using the LANG strategy further boosts the performance by 3.2%pt and 10%pt respectively as compared to the base model fine-tuned without this strategy. Surprisingly, even without the LANG strategy, models instruction-tuned on sPhinX still perform better than those tuned on other datasets with LANG.
Furthermore, employing the LANG strategy leads to additional performance improvements indicating that LANG can effectively enhance the multilingual capabilities of LLMs. From the detailed results in Appendix §A.2, we observe no performance regression on high resource languages which normally occurs due to catastrophic forgetting Chang et al. (2023).
We also observe significant performance improvements in medium and low-resource languages such as Arabic, Hindi, Thai, Turkish, Tamil, and Telugu, further showcasing the effectiveness of our dataset and the LANG fine-tuning strategy (Appendix §A.2).
6.1 Regression Analysis on Standard LLM Benchmarks
It is well-studied that training on multiple languages cause regression on the performance on English language due to catastrophic forgetting Chang et al. (2023). We test this by checking for performance of the Phi-3-small model fine-tuned with sPhinX on English in the multilingual benchmarks we evaluate ((Appendix §A.2) and on popular English-only benchmarks.
We find that the Phi-3-small sPhinX model maintains its performance in English on the multilingual benchmarks and is also consistently able to maintain performance on standard English benchmarks such as MMLU (5-shot) Hendrycks et al. (2021), MedQA (2-shot) Jin et al. (2021), Arc-C (10-shot), Arc-E (10-shot) Clark et al. (2018), PiQA (5-shot) Bisk et al. (2020), WinoGrande (5-shot) Sakaguchi et al. (2021), OpenBookQA (10-shot) Mihaylov et al. (2018), BoolQ (2-shot) Clark et al. (2019) and CommonSenseQA (10-shot) Talmor et al. (2018) (Table 10). We notice some regression in the GSM-8k (8-shot, CoT) Cobbe et al. (2021) benchmark. This indicates that gains in multilingual performance caused by sPhinX do not come at the cost of regression in English performance.
7 Conclusion
In this paper, we demonstrated how instruction tuning Phi-3-small and Mistral-7B on sPhinX effectively improve their multilingual capabilities. We observed that instruction tuning the models using the sPhinX dataset leads to consistent gains on an average by 4.2%pt for Phi-3-small and 5%pt for Mistral-7B respectively. We find that the Phi-3-small model instruction tuned on sPhinX is the best performing model on multilingual benchmarks. Moreover, sPhinX exhibits greater sample efficiency and diversity compared to other multilingual instruction tuning datasets.
Additionally, we proposed a method for further enhancing the model’s performance by utilizing LANG, which supplements the training examples with $N$ samples from a few-shot set, providing the model with additional context to aid in its learning process. This further boosts the performance for both Phi-3-small and Mistral-7B by 3.2%pt and 10%pt respectively. Models instruction-tuned on sPhinX, exhibit enhanced performance even in languages they have not previously encountered during training. Finally, we evaluate the performance of the Phi-3-small model fine-tuned on sPhinX and find that it is able to maintain English performance, suggesting that gains in multilingual performance do not come at the cost of English performance while fine-tuning with sPhinX. Through our comprehensive experiments, analyses, and findings, we aim to contribute to the progress of LLMs for multilingual purposes, promoting advancements in natural language processing across a broader spectrum of languages.
8 Future Work
We have conducted all experiments using 7B base models in full parameter fine-tuning settings. It would be interesting to study the same effect by adaptive learning using LoRA Hu et al. (2022) or PEFT Mangrulkar et al. (2022). We believe that our strategies could also be effective for models with fewer parameters, leading to a notable improvement in multilingual performance.
Our LANG strategy involves utilizing $N$ examples from the same language. Other ideas that can be explored are incorporating $N$ examples from the same language script to increase the diversity of the sample. This approach could be particularly beneficial for enhancing the performance of models in low-resource languages.
Limitations
Our study has several limitations that can be considered in future research. Firstly, we conducted an extensive series of experiments, utilizing significant GPU resources and substantial time for model fine-tuning. Due to these resource-intensive processes, it may be difficult to apply our strategies to fully fine-tune a model. Besides, our study is confined to 7B models, explicitly excluding larger models. Despite this limitation, we believe our methodologies are broadly applicable for fine-tuning smaller datasets using techniques like LoRA and PEFT. Secondly, our fine-tuning dataset focuses on reasoning tasks and excludes some low-resource languages. We evaluated the models’ performance against these reasoning benchmarks. However, we did not benchmark our models on generative tasks such as summarization, nor did we evaluate models on hallucination, toxicity, or fairness.
Ethics Statement
Despite our rigorous efforts to ensure that our dataset is free from discriminatory, biased, or false information, there remains a possibility that these problems are present, particularly in multilingual contexts. Hence, it is possible that these issues might propagate to our fine-tuned models as well. We are committed to mitigate such risks and strongly advocate for the responsible use of recipes and prevent any unintended negative consequences.
Acknowledgements
We thank the Orca team at Microsoft Research for sharing the Orca dataset to implement the synthetic dataset creation recipe.
Appendix A Appendix
A.1 Prompt Templates
Figure 6 is the template for Selective Translation that was used to generate the synthetic data. Our reference dataset is in English and the {language} is the target language to generate the data in. Figure 7, 8, 9,10 and 11 are the prompts used to evaluate XQuAD, XstoryCloze, Xwinograd, XCOPA and Belebele respectively.
A.2 Detailed Results
Tables 4, 5, 6, 7 and 8 show the granular results on our models and dataset.
[…]
Please carefully convert a conversation between a human and an AI assistant from English to language. The dialogue will be presented in JSON format, where ’system’ denotes system instructions, ’human’ indicates user queries, and ’assistant’ refers to the AI’s response. You should approach this task as if the ’human’ original language is {language}. Translate the ’system’ instructions fully into {language}. For the ’human’ input, however, carefully discern which segments require translation into {language}, while leaving other parts in their original form.
For instance: 1. If the human contains a mix of languages, only translate the instruction part.
2. If the task is about language correction do not translate the target passage.
For the ’assistant’ part, generate the ’assistant’ response as you were prompted with ths newly translated system and assistant instructions. The outcome should retain the JSON format. Your response should solely contain the JSON. Do not translate the JSON keys. {"system": System text here, "human": User text here, "assistant": Assistant text here }
Figure 6: Prompt for Selective Translation using GPT-4
[…]
The task is to solve reading comprehension problems. You will be provided questions on a set of passages and you will need to provide the answer as it appears in the passage. The answer should be in the same language as the question and the passage.
Context:
{context}
Question:
{question}
Referring to the passage above, the correct answer to the given question is {answer}
Figure 7: XQuAD evaluation prompt
[…]
{input_sentence_1} {input_sentence_2}
{input_sentence_3} {input_sentence_4}
What is a possible continuation for the story given the following options?
Option1: {sentence_quiz1} Option1: {sentence_quiz2}
Figure 8: XstoryCloze evaluation prompt
[…]
Select the correct option out of option1 and option2 that will fill in the _ in the below sentence:
{sentence}
Choices:
-option1: {option1}
-option2: {option2}
Figure 9: Xwinograd evaluation prompt
[…]
The task is to perform open-domain commonsense causal reasoning. You will be provided a premise and two alternatives, where the task is to select the alternative that more plausibly has a causal relation with the premise. Answer as concisely as possible in the same format as the examples below: Given this premise:
{premise}
What’s the best option?
-choice1 : {choice1}
-choice2 : {choice2}
We are looking for{if question == c̈ause%̈} a cause {else %} an effect {endif %}
Figure 10: XCOPA evaluation prompt
[…]
The task is to perform reading comprehension task. Given the following passage, query, and answer choices, output only the letter corresponding to the correct answer. Do not give me any explanations to your answer. Just a single letter corresponding to the correct answer will suffice.
Passage: {flores_passage}
Query: {question}
Choices:
A: {mc_answer1}
B: {mc_answer2}
C: {mc_answer3}
D: {mc_answer4}
Figure 11: Belebele evaluation prompt
Hyperparameter Value
Batch Size 512
Context length 8192
Learning Rate $10*e^{-5}$
Scheduler Cosine
Epochs 10
Weight Decay 0.1
Optimizer AdamW
Table 2: Hyperparameters for model fine-tuning
$N$ $p(N)$ $N$ $p(N)$
0 0.3 4 0.1
1 0.2 5 0.1
2 0.1 6 0.1
3 0.1
Table 3: Probabilities of selecting number of shots in the LANG strategy
Language en fr jp pt ru zh avg
Mistral-7B
Base Model 0.52 0.47 0.52 0.54 0.54 0.50 0.52
IFT 0.61 0.57 0.57 0.57 0.60 0.56 0.58
m-Alpaca 0.61 0.57 0.57 0.57 0.60 0.56 0.58
Aya 0.55 0.56 0.54 0.54 0.56 0.54 0.55
Bactrian 0.61 0.57 0.57 0.57 0.60 0.56 0.58
sPhinX-0s 0.75 0.65 0.68 0.67 0.66 0.65 0.68
sPhinX 0.80 0.69 0.72 0.70 0.67 0.67 0.71
Phi-3-small
Base Model 0.86 0.67 0.73 0.77 0.74 0.72 0.75
IFT 0.86 0.78 0.72 0.78 0.77 0.75 0.78
m-Alpaca 0.87 0.76 0.75 0.78 0.76 0.71 0.81
Aya 0.79 0.61 0.67 0.70 0.70 0.66 0.69
Bactrian 0.83 0.72 0.71 0.75 0.70 0.68 0.73
sPhinX-0s 0.89 0.77 0.78 missing0.79 0.81 0.76 0.80
sPhinX missing0.89 missing0.76 missing0.79 0.78 missing0.82 missing0.77 missing0.84
Table 4: Language-wise performance of instruction-tuned Mistral-7B and Phi-3-small  models evaluated on XWinograd (0-shot). Metric: Accuracy. The best performing IFT dataset for each model is indicated in bold, and the overall best performing IFT model is indicated with an underline.
Language ar de el en es hi ro ru th tr vi zh avg
Mistral-7B
Base Model 0.62 0.81 0.64 0.89 0.86 0.65 0.82 0.71 0.59 0.68 0.79 0.76 0.74
IFT 0.42 0.68 0.33 0.92 0.66 0.5 0.71 0.61 0.38 0.63 0.71 0.68 0.60
m-Alpaca 0.10 0.75 0.15 0.86 0.82 0.12 0.62 0.68 0.12 0.38 0.52 0.46 0.46
Aya 0.33 0.73 0.65 0.85 0.8 0.63 0.75 0.67 0.57 0.61 0.75 0.59 0.66
Bactrian 0.67 0.76 0.26 0.85 0.86 0.74 0.77 0.71 0.59 0.69 0.77 0.65 0.69
sPhinX-0s 0.54 0.76 0.7 0.88 0.84 0.69 0.77 0.66 0.52 0.64 0.71 0.60 0.69
sPhinX 0.74 0.87 0.76 0.93 0.90 0.79 0.86 0.77 0.63 0.76 0.88 0.73 0.80
Phi-3-small
Base Model 0.68 0.90 0.77 0.93 0.91 0.61 0.84 0.80 0.55 0.73 0.86 0.69 0.78
IFT 0.71 0.88 0.73 0.92 0.91 0.64 0.84 0.80 0.44 0.70 0.67 0.76 0.75
m-Alpaca 0.55 0.92 0.74 0.96 0.94 0.68 0.87 0.85 0.50 0.73 0.88 0.66 0.77
Aya 0.61 0.89 0.84 0.94 0.93 0.80 0.89 0.82 0.73 0.83 0.92 0.79 0.83
Bactrian 0.81 0.92 0.81 0.95 0.95 0.80 0.91 0.84 0.74 0.82 0.92 0.79 0.85
sPhinX-0s 0.75 0.89 0.81 0.94 0.94 0.75 0.87 0.79 0.63 0.77 0.88 0.78 0.82
sPhinX missing0.84 missing0.93 missing0.87 missing0.96 missing0.96 missing0.81 missing0.91 missing0.86 missing0.73 missing0.84 missing0.92 missing0.81 missing0.87
Table 5: Granular results for XQuAD (3-shot) on our model. Metric: F1. The best performing IFT dataset for each model is indicated in bold, and the overall best performing IFT model is indicated with an underline.
Language et ht id it qu sw ta th tr vi zh en avg
Mistral-7B
Base Model 0.54 0.51 0.72 0.81 0.49 0.52 0.50 0.53 0.58 0.62 0.78 0.93 0.63
IFT 0.52 0.52 0.69 0.79 0.50 0.51 0.50 0.54 0.57 0.63 0.75 0.90 0.62
m-Alpaca 0.51 0.50 0.52 0.63 0.50 0.50 0.50 0.51 0.51 0.49 0.65 0.74 0.55
Aya 0.57 0.54 0.64 0.67 0.53 0.56 0.57 0.62 0.56 0.61 0.64 0.78 0.61
Bactrian 0.52 0.50 0.53 0.60 0.49 0.51 0.50 0.51 0.51 0.52 0.52 0.71 0.54
sPhinX-0s 0.54 0.5 0.58 0.63 0.51 0.55 0.52 0.52 0.54 0.57 0.64 0.8 0.58
sPhinX missing0.64 0.54 0.73 0.80 0.53 0.61 0.59 0.63 0.67 0.66 0.80 0.91 0.68
Phi-3-small
Base Model 0.55 0.51 0.80 0.93 0.52 0.54 0.46 0.56 0.61 0.66 0.86 0.98 0.64
IFT 0.55 0.57 0.81 0.93 0.53 0.58 0.48 0.60 0.62 0.69 0.88 0.96 0.68
m-Alpaca 0.53 0.54 0.80 0.92 0.49 0.54 0.51 0.59 0.64 0.68 0.87 0.99 0.68
Aya 0.60 0.55 0.72 0.83 0.52 0.55 0.52 0.62 0.59 0.69 0.75 0.89 0.65
Bactrian 0.62 0.56 0.83 0.91 0.54 0.62 0.52 0.66 0.65 0.71 0.86 0.99 0.71
sPhinX-0s 0.59 0.58 0.85 0.94 0.50 0.60 0.54 0.63 0.69 missing0.72 0.89 0.96 0.71
sPhinX 0.59 missing0.60 missing0.85 missing0.94 missing0.52 missing0.57 missing0.58 missing0.68 missing0.68 0.71 missing0.90 missing0.98 missing0.72
Table 6: Granular results for XCOPA (4-shot) on our model. Metric: Accuracy. The best performing IFT dataset for each model is indicated in bold, and the overall best performing IFT model is indicated with an underline.
Language ar en es eu hi id my ru sw te zh avg
Mistral-7B
Base Model 0.65 0.89 0.83 0.56 0.62 0.76 0.52 0.81 0.56 0.52 0.80 0.68
IFT 0.70 0.96 0.92 0.54 0.69 0.79 0.57 0.90 0.58 0.54 0.88 0.73
m-Alpaca 0.53 0.73 0.7 0.51 0.51 0.57 0.50 0.66 0.52 0.52 0.71 0.59
Aya 0.64 0.86 0.81 0.56 0.71 0.73 0.60 0.82 0.67 0.60 0.81 0.71
Bactrian 0.69 0.82 0.74 0.52 0.59 0.76 0.54 0.73 0.62 0.61 0.76 0.67
sPhinX-0s 0.57 0.66 0.64 0.47 0.56 0.61 0.50 0.62 0.56 0.52 0.69 0.58
sPhinX 0.83 0.96 0.94 0.57 missing0.84 0.87 missing0.61 0.91 missing0.80 missing0.69 0.94 0.81
Phi-3-small
Base Model 0.80 0.98 0.96 0.61 0.72 0.92 0.53 0.96 0.61 0.55 0.94 0.78
IFT 0.81 0.98 0.96 0.61 0.75 0.92 0.56 0.96 0.61 0.53 0.94 0.79
m-Alpaca 0.81 0.99 0.98 0.58 0.76 0.93 0.52 0.97 0.64 0.54 0.96 0.79
Aya 0.77 0.98 0.97 0.57 0.77 0.93 0.53 0.96 0.74 0.56 0.94 0.79
Bactrian 0.83 0.98 0.98 0.61 0.83 0.94 0.54 0.97 0.79 0.63 0.94 0.82
sPhinX-0s 0.84 0.98 0.97 missing0.64 0.77 0.95 0.52 0.96 0.74 0.57 0.95 0.81
sPhinX missing0.86 missing0.99 missing0.99 0.61 0.82 missing0.96 0.54 missing0.98 0.74 0.61 missing0.97 missing0.82
Table 7: Granular results for XStoryCloze (4-shot) on our model. Metric: Accuracy. The best performing IFT dataset for each model is indicated in bold, and the overall best performing IFT model is indicated with an underline.
Language ar de es en fi fr hi it jp ko ta te vi zh avg
Mistral-7B
Base Model 0.25 0.23 0.23 0.24 0.23 0.23 0.26 0.24 0.26 0.23 0.23 0.25 0.26 0.25 0.24
IFT 0.32 0.60 0.62 0.74 0.36 0.62 0.32 0.61 0.43 0.47 0.27 0.27 0.39 0.58 0.47
m-Alpaca 0.32 0.50 0.53 0.56 0.45 0.51 0.27 0.51 0.40 0.41 0.26 0.26 0.33 0.48 0.41
Aya 0.34 0.43 0.43 0.48 0.38 0.47 0.35 0.44 0.4 0.36 0.27 0.25 0.37 0.42 0.38
Bactrian 0.24 0.27 0.25 0.25 0.26 0.27 0.24 0.28 0.26 0.26 0.23 0.23 0.34 0.28 0.26
sPhinX-0s 0.64 0.75 0.75 0.82 0.66 0.79 0.53 0.73 0.69 0.66 0.48 0.44 0.66 0.75 0.67
sPhinX 0.69 0.80 0.69 0.87 0.71 0.82 missing0.60 0.79 0.73 0.73 missing0.56 missing0.48 0.70 0.80 0.71
Phi-3-small
Base Model 0.54 0.87 0.85 0.92 0.58 0.86 0.41 0.86 0.70 0.58 0.26 0.30 0.62 0.82 0.65
IFT 0.63 0.89 0.88 0.93 0.63 0.88 0.48 0.88 0.77 0.68 0.32 0.32 0.68 0.85 0.70
m-Alpaca 0.65 0.92 0.90 0.94 0.74 0.91 0.54 0.90 0.80 0.70 0.47 0.45 0.72 0.84 0.75
Aya 0.58 0.86 0.85 0.91 0.65 0.87 0.50 0.86 0.76 0.67 0.37 0.35 0.69 0.84 0.70
Bactrian 0.67 0.88 0.88 0.92 0.70 0.88 0.51 0.86 0.77 0.70 0.37 0.37 0.74 0.86 0.72
sPhinX-0s 0.73 0.91 0.90 0.93 0.75 0.92 0.57 0.91 0.82 missing0.82 0.45 0.40 0.76 0.89 0.77
sPhinX missing0.74 missing0.93 missing0.91 missing0.94 missing0.77 missing0.93 0.58 missing0.92 missing0.84 0.76 missing0.46 0.40 missing0.78 missing0.89 missing0.79
Table 8: Granular results for Belebele (0-shot) on our model. Metric: Accuracy. The best performing IFT dataset for each model is indicated in bold, and the overall best performing IFT model is indicated with an underline.
Model XC XS XW XQ BL
Mistral-7B
Base Model 0.63 0.68 0.52 0.74 0.24
IFT 0.62 0.73 0.54 0.60 0.47
m-Alpaca 0.55 0.59 0.51 0.46 0.41
Aya 0.68 0.71 0.54 0.66 0.38
Bactrian 0.54 0.67 0.54 0.69 0.26
sPhinX-0s 0.58 0.58 0.68 0.69 0.67
sPhinX 0.68 0.81 0.71 0.80 0.71
Phi-3-small
Base Model 0.64 0.78 0.75 0.78 0.65
IFT 0.68 0.79 0.78 0.75 0.70
m-Alpaca 0.68 0.79 0.81 0.77 0.75
Aya 0.65 0.79 0.69 0.83 0.72
Bactrian 0.71 0.82 0.73 0.85 0.77
sPhinX-0s 0.71 0.81 0.80 0.82 0.79
sPhinX 0.72 0.84 0.87 0.87 0.79
Table 9: Performance of Mistral-7B and Phi-3-small instruction-tuned for 10000 training steps on various datasets. Abbreviations: XC - XCOPA, XS - XStoryCloze, XW - XWinograd, XQ - XQuAD, BL - Belebele. The best performing IFT dataset for each model is indicated in bold, and the overall best performing IFT model is indicated with an underline.
Benchmarks Base Model sPhinX
MMLU
(5-shot)
0.76 0.75
HellaSwag
(5-shot)
0.81 0.83
GSM-8k
(8-shot, CoT)
0.85 0.77
MedQA
(2-shot)
0.64 0.66
Arc-C
(10-shot)
0.90 0.90
Arc-E
(10-shot)
0.97 0.97
PIQA
(5-shot)
0.84 0.89
WinoGrande
(5-shot)
0.77 0.82
OpenBookQA
(10-shot)
0.86 0.88
BoolQ
(2-shot)
0.82 0.87
CommonSenseQA
(10-shot)
0.80 0.81
Table 10: Performance of the Phi-3-small base model and the sPhinX tuned model on standard English LLM benchmarks.
High-Resource (100k)
Spanish, Chinese Simplified, Japanese
French, German, Portuguese, Italian
Mid-Resource (50k)
Dutch, Swedish, Danish
Finnish, Russian, Norwegian
Korean, Chinese Traditional, Polish
Turkish, Arabic, Hebrew
Portuguese, Czech, Hungarian
Low-Resource (25k)
Indonesian, Thai, Greek
Slovak, Vietnamese, Slovenian
Croatian, Romanian, Lithuanian
Bulgarian, Serbian, Latvian
Ukranian, Estonian, Hindi
Burmese, Bengali, Afrikaan
Punjabi, Welsh, Icelandic
Marathi, Swahili, Nepali
Urdu, Telugu, Malayalam
Russian, Tamil, Oriya
Table 11: Language distribution and samples across three tiers
Code Languages Script Data
af Afrikaan Latin 20206
ar Arabic Arabic 26803
bn Bengali Bengal 20165
bg Bulgarian Cyrillic 17300
my Burmese Burmese 12123
zh-Hans Chinese_Simplified Han 100650
zh-Hant Chinese_Traditional Hant 32363
hr Croatian Latin 17340
cs Czech Latin 32711
da Danish Latin 36348
nl Dutch Latin 36586
en English Latin 199900
et Estonian Latin 17207
fi Finnish Latin 33622
fr French Latin 100337
de German Latin 100265
el Greek Greek 17317
he Hebrew Hebrew 24483
hi Hindi Devanagari 20240
hu Hungarian Latin 31999
is Icelandic Latin 20164
id Indonesian Latin 17297
it Italian Latin 85175
jp Japanese Japanese 98366
ko Korean Hangul 30890
lv Latvian Latin 17247
lt Lithuanian Latin 17232
ml Malayalam Malayalam 19817
mr Marathi Devanagari 20069
ne Nepali Devanagari 20092
nb Norwegian Latin 36811
or Oriya Oriya 19153
pl Polish Latin 34711
pt Portuguese Latin 37229
pa Punjabi Gurmukhi 20026
ro Romanian Latin 17149
ru Russian Cyrillic 20108
sr Serbian Latin 17165
sk Slovak Latin 17255
sl Slovenian Latin 17300
es Spanish Latin 100351
sw Swahili Latin 20170
sv Swedish Latin 36533
ta Tamil Tamil 19807
te Telugu Telugu 19947
th Thai Thai 17322
tr Turkish Latin 34405
uk Ukrainian Cyrillic 17282
ur Urdu Perso-Arabic 20162
vi Vietnamese Latin 17358
cy Welsh Latin 20207
Table 12: Language Distribution in Sphinx Dataset