### v4

# sPhinX: Sample Efficient Multilingual Instruction Fine-Tuning Through N-shot Guided Prompting

Abstract
Despite the remarkable success of large language models (LLMs) in English, a significant performance gap remains in non-English languages. To address this, we introduce a novel approach for strategically constructing a multilingual synthetic instruction tuning dataset, sPhinX . Unlike prior methods that directly translate fixed instruction-response pairs, sPhinX enhances diversity by selectively augmenting English instruction-response pairs with multilingual translations. Additionally, we propose LANGIT, a novel N-shot guided fine-tuning strategy, which further enhances model performance by incorporating contextually relevant examples in each training sample. Our ablation study shows that our approach enhances the multilingual capabilities of Mistral-7B and Phi-3-small  improving performance by an average of 39.8% and 11.2%, respectively, across multilingual benchmarks in reasoning, question answering, reading comprehension, and machine translation. Moreover, sPhinX maintains strong performance on English LLM benchmarks while exhibiting minimal to no catastrophic forgetting, even when trained on 51 languages.
1 Introduction
Large Language Models (LLMs) have demonstrated exceptional performance across various tasks in English. However, their performance in some non-English languages remains comparatively limited Ahuja et al. (2023); Asai et al. (2024). Further, the gap between the performance of Large Language Models (LLMs) and Small Language Models (SLMs) is more pronounced (Ahuja et al., 2024) in non-English languages. Cui et al. (2023) and Balachandran (2023) utilize the method of fine-tuning models on datasets focused on particular languages. However, this can lead to catastrophic forgetting, which may negatively impact performance in English Zhao et al. (2024); Aggarwal et al. (2024). Few techniques have been proposed to bridge this gap, such as incorporating better pre-training data in various languages and improving base tokenizers Xu et al. (2024); Dagan et al. (2024). However, most of these changes need to be implemented in the pre-training stage, which demands extensive data and computational resources, making it practically unfeasible in many scenarios Brown et al. (2020). Consequently, the most well-studied technique involves fine-tuning models for specific languages and tasks. Instruction fine-tuning (IFT) has become a popular technique to enhance the performance of language models in specific languages. This method combines the benefits of both the pre-training, fine-tuning, and prompting paradigms Wei et al. (2021).
Sample diversity is essential for effective instruction tuning in multilingual datasets. Many recent datasets have been generated by translating English content into other languages or by employing self-instruct techniques based on seed prompts Li et al. (2023); Taori et al. (2023). However, both methods can limit diversity. Machine translation may result in the loss of semantic nuance (Baroni and Bernardini, 2006), while self-instruct approaches often yield repetitive and homogeneous samples Wang et al. (2022). This highlights the critical need for datasets that encompass a wide range of diverse samples.
In this paper, we present a novel recipe for creating a multilingual synthetic instruction tuning dataset, sPhinX.  It comprises 1.8M instruction-response pairs in 51 languages, derived by augmenting the Orca instruction tuning dataset samplesMukherjee et al. (2023) through Selective Translated Augmentation using GPT-4 Achiam et al. (2023). We assess the effectiveness of sPhinX by fine-tuning two models — Phi-3-small and Mistral-7B — across a range of evaluation benchmarks that test various language model capabilities across discriminative and generative tasks. We compare models fine-tuned on sPhinX with those using other synthetic multilingual instruction tuning datasets like Aya Üstün et al. (2024), Multilingual Alpaca Taori et al. (2023), and Bactrian Li et al. (2023) and observe significant performance gains across languages. We also compare our proposed translation strategy with translating the entire instruction using Azure Translator API, as is done with the popular multilingual synthetic IFT datasets to demonstrate the efficacy of our approach.
The contributions of this paper are as follows:
  • We introduce a novel approach to generate synthetic data for multilingual instruction tuning by Selective Translated Augmentation of the Orca dataset with the assistance of GPT-4 (§3.1)
  • We devise LAnguage-Specific N-shot Guided Instruction Tuning (LANGIT) strategy for enhancing the multilingual capabilities of LLMs (§4)
  • We also conduct extensive instruction tuning experiments on various multilingual instruction tuning datasets to evaluate generalizability in multilingual settings (§6).
  • We plan to release a subset of the augmented dataset by applying our strategy to the OpenOrca  [^1] https://huggingface.co/datasets/Open-Orca/OpenOrca dataset (Lian et al., 2023) (Open-sPhinX) as well.
2 Related Work
2.1 Multilingual Instruction fine-tuning
Early studies focused on fine-tuning pre-trained models on a variety of languages through data augmentation for a single task Hu et al. (2020); Longpre et al. (2021); Asai et al. (2022). Currently, the approach has shifted to fine-tuning these models using a wide variety of tasks Longpre et al. (2023); Ouyang et al. (2022). Models such as BLOOMZ Muennighoff et al. (2022) and mT0 Muennighoff et al. (2022) make significant strides in improving the multilingual performance of decoder-based models Ahuja et al. (2023). There have been multiple multilingual instruction datasets and models proposed such as Bactrian (Li et al., 2023), AYA (Üstün et al., 2024), polyLM (Wei et al., 2023b) after BLOOMZ and mT0 Muennighoff et al. (2023). However, these models still do not perform as well as English in other languages, with the gap being huge for low-resource languages and languages written in scripts other than the Latin script  Ruder et al. (2021); Ahuja et al. (2023); Asai et al. (2024); Ahuja et al. (2024). In this work, we aim to narrow the performance gap by introducing a strategy for creating datasets for multilingual instruction tuning and recipes for fine-tuning, which we will discuss in the following sections.
Figure 1: The figure above illustrates pipelines for sPhinX data creation using Selective Translated Augmentation and Multilingual Instruction Tuning using LANGIT strategy.
Figure 2: The figure compares the Translation API with Selective Translated Augmentation. The Translation API translates the entire input into Russian, while the Selective strategy localizes only necessary components. Here, system and user prompts are translated, but the input question and assistant’s response remain in the original language, preserving structure and intent.
2.2 Multilingual Synthetic Data Generation
Most instruction-tuning datasets across multiple languages typically focus on general tasks rather than specific reasoning capabilities. Although datasets like Orca Mukherjee et al. (2023) and Orca 2 Mitra et al. (2023) exist in English, they highlight a prevalent issue: current methods often prioritize style imitation over leveraging the reasoning abilities found in large foundation models (LFMs). The Orca dataset addresses this by imitating rich signals from GPT-4, including explanation traces and step-by-step thought processes Wei et al. (2023a), guided by assistance from ChatGPT. In order to create multilingual datasets, researchers commonly use translation APIs or LLMs to translate English-specific datasets into target languages. For example, the Bactrian dataset Li et al. (2023) translates Alpaca and Dolly instructions into 52 languages using the Google Translator API and generates outputs with GPT-3.5 turbo. Another example is of this work  Lai et al. (2024), which also utilizes Google Translation API for translating their source datasets. Our dataset approach aims to tackle these challenges by selectively translating only the essential portions of multilingual inputs. This strategy not only preserves semantic information but also accommodates diverse linguistic contexts, thereby enhancing the overall quality and applicability of instruction-tuning datasets across languages. In our work, we also show the pitfalls of training with data generated only using the Translation APIs such Google Translate or Azure Translate.
3 sPhinX Dataset
In this section, we describe our dataset construction methodology (§3.1), dataset filtering, and cleaning pipelines (§3.2).
3.1 Dataset Construction
Inspired by (Mukherjee et al., 2023)’s work, we utilized the 1M GPT-4 generated instruction-response pairs from Orca and constructed our own dataset along similar lines using Selective Translated Augmentation into 50 different languages with the help of GPT-4222GPT-4 inference hyper-parameters in Azure OpenAI interface set as: temperature=0.0. We categorize them into three groups: high-resource, mid-resource, and low-resource languages as outlined in Table 7. For high-resource languages, we randomly sample 100k instruction-response pairs from the Orca 1M dataset and generate the responses from GPT-4 with Selective Translated Augmentation as shown in Figure 2. Similarly, we leverage the same strategy for medium and low-resource languages by sampling 50k and 25k pairs respectively. Although GPT-4 performs competitively with commercial translation systems (Google Translate & Bing Translate) it still lags on medium and low resource languages Jiao et al. (2023); Hendy et al. (2023). Furthermore, as highlighted in (Chang et al., 2023; Lin et al., 2023; Xia et al., 2024), fine-tuning with a large set of samples from medium and low-resource languages might lead to catastrophic forgetting of high-resource languages. Therefore, we deliberately create fewer samples for medium and low-resource languages than for high-resource ones. Besides, Shaham et al. (2024) also demonstrates that a small number of multilingual training samples is sufficient to significantly boost multilingual performance, validating our approach of using fewer samples from medium- and low-resource languages.
A fundamental problem with using an off-the-shelf translation API is the lack of semantic and task awareness, in addition to translationese Baroni and Bernardini (2006), which can result in poor quality training data. Consider for example the task of Machine Translation as part of the instruction, wherein the language of the source sentence should be retained. However, an off-the-shelf API, without task awareness, would translate it, resulting in an ambiguous instruction. To mitigate this issue, we used GPT-4 to augment the instructions using Selective Translated Augmentation, so that task-specific components of instruction responses are translated into the appropriate language without changing the semantic meaning. Figure 12 illustrates this with concrete examples. The first example demonstrates the aforementioned translation inconsistency issue for an instruction asking for a French equivalent of an English phrase. The second example demonstrates the consequence of direct translations in the m-Alpaca dataset: wherein the translation of the task input results in the task being ill-defined based on the instructions. As demonstrated, our proposed Selective Translated Augmentation method is able to keep the semantic information of the task intact while translating the instructions. For the exact prompt, please refer to Figure 4 in the Appendix.
3.2 Dataset Filtering and Quality Assessment
After creating the dataset, we filtered out samples where GPT-4 failed to generate a response. The final dataset comprised 1.8 million samples in 51 languages(Table: 15), divided into three subsets: Train, Validation, and Few-shot. Each language’s dataset was partitioned to ensure that the Validation and Few-shot sets contained 2,000 and 1,000 samples, respectively, while the Train set included the remaining data. This approach guarantees consistent distributions across languages in the Validation and Few-shot sets, ensuring equitable representation regardless of the training distribution. The final split ratio for Train, Validation, and Few-shot sets was 92:5.3:2.7.
We also conducted a small-scale quality assessment of the generated data for languages such as Bengali, Hindi, German, Turkish, and Tamil. The researchers and engineers in our organization, who are native speakers of these languages, evaluated the data on the basis of coherence, fluency, and information retention. Our findings indicate that the generated dataset is moderate to high quality.
3.3 Sample Diversity in sPhinX
Unlike prior multilingual datasets such as Bactrian and m-Alpaca, which translate a fixed set of instruction-response pairs into multiple languages, sPhinX ensures diversity by sampling unique subsets of instruction-response pairs for each language.
For instance, Bactrian is constructed from 67k English instruction-response pairs (Alpaca + Dolly) and translated into 52 languages, resulting in identical samples across all languages. In contrast, sPhinX samples from 1M GPT-4-generated instruction-response pairs, ensuring that no two languages share the exact same subset.
Mathematically, the probability that all samples in one language dataset $A$ are also in another language dataset $B$, when sampled without replacement from a larger dataset $D$, is given by:
$P(\text{all A in B})\approx\left(\frac{m}{N}\right)^{n}=\left(\frac{100{,}000}{1{,}000{,}000}\right)^{20{,}000}$
where:
  • $N=1{,}000{,}000$ (Total samples in sPhinX),
  • $n=20{,}000$ (Samples in language $A$),
  • $m=100{,}000$ (Samples in language $B$).
Since the exponential term results in an extremely small probability, this confirms that no two languages have identical instruction-response sets in sPhinX.
To further enhance diversity, we apply Selective Translated Augmentation, translating 10% of samples for high-resource languages, 5% for mid-resource languages, and 2.5% for low-resource languages. This ensures that translated content varies across languages, preventing uniformity.
Additionally, code-switching naturally emerges from this augmentation process, further increasing linguistic diversity. Compared to Aya, which exhibits moderate variation across task instructions, sPhinX introduces greater sample diversity by leveraging a larger and more heterogeneous seed set (Mukherjee et al., 2023) and selective augmentation strategy. Exploring code-switching phenomena would be an interesting task in these synthetically generated datasets, but currently that is out-of-scope of for this work.
4 LANGIT
Following Instruction Tuning strategies of  (Longpre et al., 2023) and also taking inspiration from (Min et al., 2022), we devise Language-Specific N-shot Guided Instruction fine-tuning (LANGIT) Figure: 1. This method aims to improve the model’s ability to follow instructions by augmenting training examples with additional context from a set of few-shot examples in the same language. This added context helps guide the model, enabling it to learn more effectively from the provided examples.
For each training example, we begin by sampling a number of few-shot examples, which are instruction-response pairs in the same language. The number of few-shot examples $N$ is determined probabilistically, with a 30% chance of selecting no few-shot examples, a 20% chance of selecting one, and gradually lower probabilities for higher numbers of few-shots. The maximum number of few-shot examples we sample is six, due to constraints imposed by the model’s context length (8192 tokens) and the typically higher tokenization length in languages other than English.
Once the number of few-shot examples is determined, they are prepended to the main training example, forming an augmented input. This augmented input is then fed into the model for instruction tuning. The purpose of this approach is to expose the model to additional examples of different tasks, helping it generalize better to new tasks in the same language.
We performed experiments to analyze how the model performs on each dataset when fine-tuned using the LANGIT strategy detailed in the next section (§6). Additionally, we fine-tuned the models on the sPhinX dataset without using LANGIT to provide a baseline for comparison. To assess the effectiveness of each instruction-tuning dataset on an equal scale, we conducted a comparative analysis of model performance on different benchmarks, fine-tuning each model on approximately 8 billion tokens per dataset using the LANG strategy.
This fine-tuning strategy is consistently applied across datasets for both the Phi-3-small and Mistral-7B base models. A comparison of token lengths across different datasets is provided in Table 6, showing the average token lengths as tokenized by the Phi-3-small model.
5 Experiments
5.1 Setup
Base Models: We use Mistral-7B [^3] We specifically use the v1.0 base model from
https://huggingface.co/mistralai/Mistral-7B-v0.1
and Phi-3-small Abdin et al. (2024) base model variants and instruction fine-tune them.
Datasets: Apart from the sPhinX dataset, we use Bactrian Li et al. (2023), m-Alpaca Wei et al. (2023b) and Aya Singh et al. (2024b) instruction datasets for comparative evaluation. We also utilize the Azure Translator API [^4] https://azure.microsoft.com/en-us/products/ai-services/ai-translator (sPhinX-T) to translate the original dataset into all our target languages, demonstrating the effectiveness of our Selective Translated Augmentation approach. More details about the datasets used for comparative evaluation are present in Appendix §A.2.
Evaluation: We evaluate [^5] Evaluation prompts and other details in Appendix §A.1 and §A.3. our fine-tuned models along with the available base and Instruction fine-tuned model variants of Mistral-7B and Phi-3-small (IFT [^6] We take the Mistral-7B instruction-tuned variant from https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.1 and Phi-3-small variant from https://huggingface.co/microsoft/Phi-3-small-8k-instruct.) on 4 discriminative tasks; XCOPA Ponti et al. (2020)(4-shot), XStoryCloze Lin et al. (2022)(4-shot), XWinograd Muennighoff et al. (2023)(0-shot), Tikhonov and Ryabinin (2021), Belebele(0-shot) Bandarkar et al. (2023), and 2 generative tasks; XQuAD (3-shot) Artetxe et al. (2020) and Translation (4-shot) Bojar et al. (2014, 2016); Kocmi et al. (2023) using the language model evaluation harness Gao et al. (2023). The number of few-shot selection are inspired from these works Ahuja et al. (2023, 2024); Asai et al. (2024)
Apart from generative tasks such as XQuAD, and machine translation, we also evaluate our instruction-tuned models on open-ended generation prompts. For this, we use an LLM-based evaluation approach to simulate win rates. We use the open-source test set from the Aya Dataset (Singh et al., 2024a), which includes 250 prompts per language across six languages. We use GPT-4o as the LLM evaluator to pick the preferred model generation on this test set, and we subsequently compute win rates (%) based on these preferences. To avoid a potential bias, we randomize the order of the models during the evaluation. The prompt for the evaluator is described in the Appendix: §A.1.
6 Results
Figure 3: Performance of Mistral-7B and Phi-3-small when instruction-tuned on 8B tokens across various datasets on different benchmarks. The solid lines represent the Phi fine-tuned models and the dashed lines represent the Mistral fine-tuned models.
Model $XC$ $XS$ $XW$ $XQ$ $BL$ $MT^{1}$ $MT^{2}$
Mistral-7B
Base Model 0.63 0.68 0.52 0.74 0.24 0.54 0.42
IFT 0.62 0.73 0.54 0.60 0.47 0.49 0.39
m-Alpaca 0.55 0.59 0.51 0.46 0.41 0.41 0.39
Aya 0.68 0.71 0.54 0.66 0.38 0.39 0.37
Bactrian 0.54 0.67 0.54 0.69 0.26 0.45 0.34
sPhinX-T 0.61 0.78 0.57 0.78 0.67 0.49 0.38
sPhinX-0s 0.58 0.58 0.68 0.69 0.67 0.49 0.42
sPhinX 0.68 0.81 0.71 0.80 0.71 0.55 0.46
Phi-3-small
Base Model 0.64 0.78 0.75 0.78 0.65 0.54 0.42
IFT 0.68 0.79 0.78 0.75 0.70 0.54 0.46
m-Alpaca 0.68 0.79 0.81 0.77 0.75 0.45 0.39
Aya 0.65 0.79 0.69 0.83 0.72 0.41 0.40
Bactrian 0.71 0.82 0.73 0.85 0.77 0.54 0.40
sPhinX-T 0.70 0.80 0.77 0.78 0.75 0.55 0.44
sPhinX-0s 0.71 0.81 0.80 0.82 0.79 0.56 0.45
sPhinX 0.72 0.84 0.87 0.87 0.79 0.56 0.46
Table 1: Performance of Mistral-7B and Phi-3-small instruction-tuned on various datasets. Abbreviations: $XC$ XCOPA (Acc.,4-shot), $XS$ XStoryCloze (Acc.,4-shot) , $XW$ XWinograd (Acc., 0-shot), $XQ$ XQuAD (F1,3-shot), $BL$ Belebele (Acc., 0-shot). $MT^{1}$- Translation for x:en (ChrF, 4-shot), $MT^{2}$ Translation for en:x direction (ChrF, 4-shot). The best performing dataset for each model is indicated in bold, and the overall best performing model is indicated with an underline.
We evaluate reasoning, question answering, translation and reading comprehension abilities of the Phi-3-small and Mistral-7B models, instruction-tuned on different multilingual datasets, using various benchmarks and find that fine-tuning on sPhinX provides an average improvement of 39.8% and 11.2% respectively on both the models. (Refer to sPhinX-0s in Table 1 for overall results and to Appendix §A.5 for language-wise results). Additionally, as observed in Figure 3, the sPhinX dataset significantly enhances the multilingual performance of the Phi-3-small and Mistral-7B model compared to other datasets even when fine-tuned on an equal number of tokens.
7 Ablations
7.1 Improvements from LANGIT
To demonstrate the effectiveness of our LANGIT strategy, we also instruction-tuned the models on sPhinX with 0 shots, referring to this as sPhinX-0s. As shown in Table 1 (with detailed results in Appendix §A.5), models fine-tuned on sPhinX especially Mistral-7B exhibit superior performance compared to its counterparts fine-tuned on other datasets across all benchmarks. Moreover, fine-tuning both Mistral-7B and the Phi-3-small on sPhinX using the LANGIT strategy further boosts the performance by an average of 15% and 3.2% respectively as compared to the vanilla fine-tuned model (sPhinX -0s) across multilingual benchmarks.
Furthermore, employing the LANGIT strategy leads to additional performance improvements indicating that LANGIT can effectively enhance the multilingual capabilities of LLMs. From the detailed results in Appendix §A.5, we observe no performance regression on high resource languages which normally occurs due to catastrophic forgetting Chang et al. (2023).
We also observe significant performance improvements in medium and low-resource languages such as Arabic, Hindi, Thai, Turkish, Tamil, and Telugu, further showcasing the effectiveness of our dataset and the LANGIT fine-tuning strategy (Appendix §A.5).
7.2 Comparisons with the API translated dataset
Due to the code-mixed nature of the instruction along with CoT reasoning explanations, a single sample of sPhinX is notably richer as compared to its counterparts from the other datasets. This can be observed in the Table 1 for sPhinX-T wherein the sPhinX trained models with the LANGIT strategy outperform the directly translated dataset baselines by an average of 11.7% and 6.3% for both Mistral-7B and Phi-3-small respectively when compared to the sPhinX-T baselines across multilingual benchmarks. Consequently, even with fewer samples as compared to the other datasets (keeping the number of the tokens the same), models trained on sPhinX achieve better performance, thereby demonstrating the per-sample efficiency of sPhinX.
7.3 Simulated Preference Evaluation
As shown in Table 2, our win-rate experiments reveal that the GPT-4o evaluator predominantly favored outputs generated by the Mistral base model trained on the sPhinX dataset using the LANGIT strategy over other models. For the Phi baselines, we observed a higher percentage of TIEs for all the languages except English, where the evaluator rated both outputs equally, rather than favoring a specific model. This performance gap between the Mistral and Phi models likely arises from the age of their respective base models. Since the Mistral base model is older, it benefits more from additional training on our dataset, whereas the more recently released Phi models are already competitive enough on these benchmarks resulting in preferring both the outputs equally.
Model ar en po te tu zh
Mistral-7B
IFT 75 / 9 / 16 59 / 36 / 5 57 / 27 / 16 62 / 15 / 23 65 / 13 / 22 70 / 26 / 4
m-Alpaca 85 / 2 / 13 74 / 21 / 5 52 / 33 / 15 69 / 8 / 23 70 / 4 / 25 75 / 15 / 10
Aya 78 / 11 / 11 85 / 11 / 4 55 / 30 / 15 56 / 18 / 25 62 / 19 / 19 78 / 14 / 8
Bactrian 82 / 4 / 13 85 / 12 / 3 57 / 31 / 12 56 / 18 / 26 62 / 20 / 18 74 / 14 / 12
sPhinX-T 61 / 23 / 16 63 / 27 / 10 56 / 24 / 20 56 / 24 / 20 62 / 19 / 19 66 / 23 / 11
sPhinX-0s 65 / 19 / 16 74 / 20 / 5 55 / 31 / 14 51 / 22 / 27 52 / 36 / 12 68 / 20 / 11
Phi-3-small
IFT 46 / 38 / 17 55 / 43 / 2 50 / 44 / 6 39 / 29 / 36 30 / 27 / 43 50 / 44 / 6
m-Alpaca 46 / 5 / 49 59 / 29 / 12 44 / 21 / 35 33 / 6 / 60 31 / 10 / 59 30 / 6 / 64
Aya 40 / 18 / 42 80 / 11 / 9 56 / 16 / 28 37 / 18 / 46 25 / 21 / 54 36 / 22 / 41
Bactrian 32 / 13 / 55 68 / 17 / 15 51 / 14 / 36 24 / 14 / 62 32 / 15 / 53 26 / 11 / 63
sPhinX-T 35 / 14 / 51 75 / 12 / 13 61 / 12 / 27 23 / 14 / 63 36 / 18 / 47 37 / 12 / 51
sPhinX-0s 23 / 11 / 66 61 / 17 / 22 32 / 18 / 50 14 / 9 / 77 15 / 10 / 74 14 / 7 / 79
Table 2: Win rates (%) according to GPT-4o: The first value represents the percentage of outputs where the evaluator preferred the sPhinX and LANGIT trained model. The second value indicates the percentage of outputs preferred from the target model. The third value reflects cases where the evaluator rated both outputs equally (TIE).
Benchmarks Base Model sPhinX
MMLU
(5-shot)
0.76 0.75
HellaSwag
(5-shot)
0.81 0.83
GSM-8k
(8-shot, CoT)
0.85 0.77
MedQA
(2-shot)
0.64 0.66
Arc-C
(10-shot)
0.90 0.90
Arc-E
(10-shot)
0.97 0.97
PIQA
(5-shot)
0.84 0.89
WinoGrande
(5-shot)
0.77 0.82
OpenBookQA
(10-shot)
0.86 0.88
BoolQ
(2-shot)
0.82 0.87
CommonSenseQA
(10-shot)
0.80 0.81
Table 3: Performance of the Phi-3-small base model and the sPhinX tuned model on standard English LLM benchmarks.
7.4 Regression Analysis on Standard LLM Benchmarks
It is well studied that training in multiple languages causes regression in performance in English due to catastrophic forgetting Chang et al. (2023). We test this phenomenon for our trained models by checking the performance of the Phi-3-small model fine-tuned with sPhinX on English in the multilingual benchmarks we evaluate ((Appendix §A.5) and on popular English-only benchmarks (Table 3).
We find that the Phi-3-small fine-tuned on sPhinX maintains its performance in English on the multilingual benchmarks and is also consistently able to maintain performance on standard English benchmarks such as MMLU (5-shot) Hendrycks et al. (2021), MedQA (2-shot) Jin et al. (2021), Arc-C (10-shot), Arc-E (10-shot) Clark et al. (2018), PiQA (5-shot) Bisk et al. (2020), WinoGrande (5-shot) Sakaguchi et al. (2021), OpenBookQA (10-shot) Mihaylov et al. (2018), BoolQ (2-shot) Clark et al. (2019) and CommonSenseQA (10-shot) Talmor et al. (2018) (Table 3). We notice some regression in the GSM-8k (8-shot, CoT) Cobbe et al. (2021) benchmark. This indicates that gains in multilingual performance caused by sPhinX do not come at the cost of regression in English performance.
8 Conclusion
In this paper, we demonstrated how instruction tuning Mistral-7B and Phi-3-small on sPhinX effectively improve their multilingual capabilities. We observed that instruction tuning the models using the sPhinX dataset leads to performance improvement by an average of 39.8% and 11.2% for Mistral-7B and Phi-3-small respectively when compared to their corresponding base models across multilingual benchmarks. Moreover, sPhinX exhibits greater sample efficiency and diversity compared to other multilingual instruction tuning datasets. We also proposed LANGIT, a strategy that enhances model performance by incorporating $N$ few-shot examples, boosting results by 15% and 3.2% for Mistral-7B and Phi-3-small, respectively, over vanilla fine-tuning with sPhinX. Compared to the sPhinX-T translation baseline, LANGIT yielded gains of 11.7% and 6.3%. Models fine-tuned on sPhinX also showed improved performance in unseen languages without degrading English performance. We also observed that the GPT-4o win-rate evaluations favored Mistral-7B with LANGIT, while Phi-3-small showed more ties due to its stronger baseline. Finally, we plan on releasing a subset of our augmented dataset built on OpenOrca (Open-sPhinX).
9 Future Work
All experiments were conducted using 7B base models with full fine-tuning. It would be interesting to explore our methods with adaptive fine-tuning techniques like LoRA Hu et al. (2022) or PEFT Mangrulkar et al. (2022), and on smaller models, where we expect similar gains in multilingual performance. Our LANGIT strategy uses $N$ examples from the same language and future work could investigate using $N$ examples from the same script to introduce greater diversity, especially for improving performance in low-resource languages. We also observed code-switched data being generated when we employed this strategy to generate data. It will be interesting to explore this phenomena in a future study.
Limitations
Our study has several limitations that can be considered in future research. Firstly, we conducted an extensive series of experiments, utilizing significant GPU resources and substantial time for model fine-tuning. Due to these resource-intensive processes, it may be difficult to apply our strategies to fully fine-tune a model. Besides, our study is confined to 7B models, explicitly excluding larger models. Despite this limitation, we believe that our methodologies are broadly applicable for fine-tuning smaller datasets using techniques like LoRA and PEFT. While some languages were not included, we made a conscious effort to cover a diverse set spanning multiple scripts and language families.
Ethics Statement
Despite our rigorous efforts to ensure that our dataset is free from discriminatory, biased, or false information, there remains a possibility that these problems are present, particularly in multilingual contexts. Hence, it is possible that these issues might propagate to our fine-tuned models as well. We are committed to mitigating such risks and strongly advocate for the responsible use of recipes and prevent any unintended negative consequences.
Appendix A Appendix
A.1 Prompt Templates
Figure 4 is the template for Selective Translated Augmentation that was used to generate the synthetic data. Our reference dataset is in English and the {language} is the target language to generate the data in. Figure 5, 6, 7,8, 9 and 10 are the prompts used to evaluate XQuAD, XstoryCloze, Xwinograd, XCOPA, Belebele and Translation respectively. Figure 11 denotes the prompt used for simulating win-rate evaluations.
A.2 Baseline Datasets For Comparative Evaluation
  • Bactrian Li et al. (2023) is a machine translated dataset of the original alpaca-52k Taori et al. (2023) and dolly-15k Conover et al. (2023) datasets into 52 languages. The instructions for this dataset were translated using a Translation API and then GPT-3.5-Turbo was prompted to generate outputs. We fine-tune our models on the complete dataset consisting of 3.4M instances.
  • m-Alpaca Wei et al. (2023b) is a self-instruct dataset that translates seed instructions from English to 11 languages, using GPT-3.5-Turbo for response generation. We fine-tune our models on the full dataset, which contains 500k data points.
  • Aya Singh et al. (2024a) contains human-curated prompt-completion pairs in 65 languages, along with 44 monolingual and multilingual instruction datasets and 19 translated datasets across 114 languages, totaling around 513M instances. To ensure parity with the sPhinX dataset, we sampled it down to 2.7M instances, ensuring equal representation for each language in our subset.
A.3 Evaluation Benchmarks
  • XCOPA: A causal commonsense reasoning dataset in 11 languages, evaluated in a 4-shot prompt setting.
  • XStoryCloze: A professionally translated version of the English StoryCloze dataset Mostafazadeh et al. (2017) in 10 languages, evaluated in a 4-shot prompt setting.
  • Belebele: A parallel reading comprehension dataset across 122 languages, with evaluation on a subset of 14 languages in a 0-shot prompt setting.
  • XQuAD: A QA dataset consisting of professional translations of a subset of SQuAD into 10 languages, evaluated in a 3-shot prompt setting due to context window limitations.
  • XWinograd: A collection of Winograd Schemas in six languages for cross-lingual commonsense reasoning, evaluated in a 0-shot setting.
  • Translation: We utilize a subset of WMT14, WMT16 and WMT23 of language pairs (7 languages), with evaluation in a 4-shot setting.
A.4 Hyperparameters and Training Setup
We used 5 nodes with each node containing 8 A100 GPUs with 80GB VRAM. These nodes communicated with each other using InfiniBand  [^7] https://network.nvidia.com/pdf/whitepapers/IB_Intro_WP_190.pdf. We use DeepSpeed Rasley et al. (2020) to do distributed fine-tuning over these GPUs. We use the same hyperparameters (Table 4) to fine-tune both Mistral-7B and Phi-3-small models.
A.5 Detailed Results
Tables 8, 9, 10, 11, 12, 13 and 14 show the granular results on our models and dataset.
[…]
Please carefully convert a conversation between a human and an AI assistant from English to language. The dialogue will be presented in JSON format, where ’system’ denotes system instructions, ’human’ indicates user queries, and ’assistant’ refers to the AI’s response. You should approach this task as if the ’human’ original language is {language}. Translate the ’system’ instructions fully into {language}. For the ’human’ input, however, carefully discern which segments require translation into {language}, while leaving other parts in their original form.
For instance: 1. If the human contains a mix of languages, only translate the instruction part.
2. If the task is about language correction do not translate the target passage.
For the ’assistant’ part, generate the ’assistant’ response as you were prompted with ths newly translated system and assistant instructions. The outcome should retain the JSON format. Your response should solely contain the JSON. Do not translate the JSON keys. {"system": System text here, "human": User text here, "assistant": Assistant text here }
Figure 4: Prompt for Selective Translation using GPT-4
[…]
The task is to solve reading comprehension problems. You will be provided questions on a set of passages and you will need to provide the answer as it appears in the passage. The answer should be in the same language as the question and the passage.
Context:
{context}
Question:
{question}
Referring to the passage above, the correct answer to the given question is {answer}
Figure 5: XQuAD evaluation prompt
[…]
{input_sentence_1} {input_sentence_2}
{input_sentence_3} {input_sentence_4}
What is a possible continuation for the story given the following options?
Option1: {sentence_quiz1} Option1: {sentence_quiz2}
Figure 6: XstoryCloze evaluation prompt
[…]
Select the correct option out of option1 and option2 that will fill in the _ in the below sentence:
{sentence}
Choices:
-option1: {option1}
-option2: {option2}
Figure 7: Xwinograd evaluation prompt
[…]
The task is to perform open-domain commonsense causal reasoning. You will be provided a premise and two alternatives, where the task is to select the alternative that more plausibly has a causal relation with the premise. Answer as concisely as possible in the same format as the examples below: Given this premise:
{premise}
What’s the best option?
-choice1 : {choice1}
-choice2 : {choice2}
We are looking for{if question == c̈ause%̈} a cause {else %} an effect {endif %}
Figure 8: XCOPA evaluation prompt
[…]
The task is to perform reading comprehension task. Given the following passage, query, and answer choices, output only the letter corresponding to the correct answer. Do not give me any explanations to your answer. Just a single letter corresponding to the correct answer will suffice.
Passage: {flores_passage}
Query: {question}
Choices:
A: {mc_answer1}
B: {mc_answer2}
C: {mc_answer3}
D: {mc_answer4}
Figure 9: Belebele evaluation prompt
[…]
Translate the following sentence pairs:
{Source Language}: {Source Phrase} {Target Language}: {Target Phrase}
Figure 10: Translation evaluation prompt
[…]
System: You are a helpful following assistant whose goal is to select the preferred (least wrong) output for a given instruction in {LANGUAGE_NAME}.
User: Which of the following answers is the best one for given instruction in {LANGUAGE_NAME}. A good answer should follow these rules: 1) It should be in {[LANGUAGE_NAME}.
2) It should answer the request in the instruction.
3) It should be factually and semantically comprehensible.
4) It should be grammatically correct and fluent.
Instruction: {INSTRUCTION}
Answer (A): {COMPLETION A}
Answer (B): {COMPLETION B}
FIRST provide a one-sentence comparison of the two answers, explaining which you prefer and why. SECOND, on a new line, state only ‘Answer (A)’ or ‘Answer (B)’ to indicate your choice. If the both answers are equally good or bad, state ‘TIE’. Your response should use the format:
Comparison: <one-sentence comparison and explanation>
Preferred: <‘Answer (A)’ or ‘Answer (B)’ or ‘TIE’>
Figure 11: Preference simulation prompt taken from (Üstün et al., 2024) evaluation suite to evaluate our models on free-form generation using GPT-4o.
Hyperparameter Value
Batch Size 512
Context length 8192
Learning Rate $10^{-5}$
Scheduler Cosine
Epochs 10
Weight Decay 0.1
Optimizer AdamW
Table 4: Hyperparameters for model fine-tuning
$N$ $p(N)$ $N$ $p(N)$
0 0.3 4 0.1
1 0.2 5 0.1
2 0.1 6 0.1
3 0.1
Table 5: Probabilities of selecting number of shots in the LANG strategy
Dataset Average Token
Length/Sample
Aya 2240
Bactrian 2465
m-Alpaca 1620
sPhinX-0s 544
sPhinX 3100
Table 6: Average Token Length in each dataset
High-Resource (100k)
Spanish, Chinese Simplified, Japanese
French, German, Portuguese, Italian
Mid-Resource (50k)
Dutch, Swedish, Danish
Finnish, Russian, Norwegian
Korean, Chinese Traditional, Polish
Turkish, Arabic, Hebrew
Portuguese, Czech, Hungarian
Low-Resource (25k)
Indonesian, Thai, Greek
Slovak, Vietnamese, Slovenian
Croatian, Romanian, Lithuanian
Bulgarian, Serbian, Latvian
Ukranian, Estonian, Hindi
Burmese, Bengali, Afrikaan
Punjabi, Welsh, Icelandic
Marathi, Swahili, Nepali
Urdu, Telugu, Malayalam
Russian, Tamil, Oriya
Table 7: Language distribution and samples across three tiers
Figure 12: Some examples of input queries and its counterpart existing in the hindi version of the Multialpaca dataset and if it was generated using the Selective Translated Augmentation strategy. Again, we observe that the samples generated using Selective Translated Augmentation translate only the required amount of information as controlled via prompting whereas in Multialpaca the translations are direct translations where only a part of the instructions have been followed to translate the input queries.
Language en fr jp pt ru zh avg
Mistral-7B
Base Model 0.52 0.47 0.52 0.54 0.54 0.50 0.52
IFT 0.61 0.57 0.57 0.57 0.60 0.56 0.58
m-Alpaca 0.61 0.57 0.57 0.57 0.60 0.56 0.58
Aya 0.55 0.56 0.54 0.54 0.56 0.54 0.55
Bactrian 0.61 0.57 0.57 0.57 0.60 0.56 0.58
sPhinX-T 0.58 0.61 0.57 0.53 0.54 0.57 0.57
sPhinX-0s 0.75 0.65 0.68 0.67 0.66 0.65 0.68
sPhinX 0.80 0.69 0.72 0.70 0.67 0.67 0.71
Phi-3-small
Base Model 0.86 0.67 0.73 0.77 0.74 0.72 0.75
IFT 0.86 0.78 0.72 0.78 0.77 0.75 0.78
m-Alpaca 0.87 0.76 0.75 0.78 0.76 0.71 0.81
Aya 0.79 0.61 0.67 0.70 0.70 0.66 0.69
Bactrian 0.83 0.72 0.71 0.75 0.70 0.68 0.73
sPhinX-T 0.87 0.74 0.74 0.77 0.77 0.72 0.77
sPhinX-0s 0.88 0.75 0.78 missing0.79 0.81 0.76 0.80
sPhinX missing0.89 missing0.76 missing0.79 missing0.79 missing0.82 missing0.77 missing0.84
Table 8: Language-wise performance of instruction-tuned Mistral-7B and Phi-3-small  models evaluated on XWinograd (0-shot). Metric: Accuracy. The best performing IFT dataset for each model is indicated in bold, and the overall best performing IFT model is indicated with an underline.
Language ar de el en es hi ro ru th tr vi zh avg
Mistral-7B
Base Model 0.62 0.81 0.64 0.89 0.86 0.65 0.82 0.71 0.59 0.68 0.79 0.72 0.73
IFT 0.42 0.68 0.33 0.92 0.66 0.5 0.71 0.61 0.38 0.63 0.71 0.68 0.60
m-Alpaca 0.10 0.75 0.15 0.86 0.82 0.12 0.62 0.68 0.12 0.38 0.52 0.46 0.46
Aya 0.33 0.73 0.65 0.85 0.80 0.63 0.75 0.67 0.57 0.61 0.75 0.59 0.66
Bactrian 0.67 0.76 0.26 0.85 0.86 0.74 0.77 0.71 0.59 0.69 0.77 0.65 0.69
sPhinX-T 0.71 0.83 0.75 0.92 0.89 0.77 0.84 0.75 0.60 0.77 0.86 0.63 0.78
sPhinX-0s 0.54 0.76 0.70 0.88 0.84 0.69 0.77 0.66 0.52 0.64 0.71 0.60 0.69
sPhinX 0.74 0.87 0.77 0.93 0.90 0.79 0.86 0.77 0.63 0.77 0.88 0.73 0.80
Phi-3-small
Base Model 0.68 0.90 0.77 0.93 0.91 0.61 0.84 0.80 0.55 0.73 0.86 0.69 0.78
IFT 0.71 0.88 0.73 0.92 0.91 0.64 0.84 0.80 0.44 0.70 0.67 0.76 0.75
m-Alpaca 0.55 0.92 0.74 0.96 0.94 0.68 0.87 0.85 0.50 0.73 0.88 0.66 0.77
Aya 0.61 0.89 0.84 0.94 0.93 0.80 0.89 0.82 0.73 0.83 0.91 0.79 0.83
Bactrian 0.81 0.92 0.81 0.95 0.95 0.80 0.90 0.84 0.72 0.82 0.91 0.79 0.85
sPhinX-T 0.80 0.91 0.82 0.95 0.94 0.80 0.90 0.83 0.69 0.81 0.91 0.73 0.78
sPhinX-0s 0.75 0.89 0.81 0.94 0.94 0.75 0.87 0.79 0.63 0.77 0.88 0.78 0.82
sPhinX missing0.84 missing0.93 missing0.87 missing0.96 missing0.96 missing0.81 missing0.91 missing0.86 missing0.73 missing0.84 missing0.92 missing0.81 missing0.87
Table 9: Granular results for XQuAD (3-shot) on our model. Metric: F1. The best performing IFT dataset for each model is indicated in bold, and the overall best performing IFT model is indicated with an underline.
Language et ht id it qu sw ta th tr vi zh en avg
Mistral-7B
Base Model 0.54 0.51 0.72 0.81 0.49 0.52 0.50 0.53 0.58 0.62 0.78 0.93 0.63
IFT 0.52 0.52 0.69 0.79 0.50 0.51 0.50 0.54 0.57 0.63 0.75 0.90 0.62
m-Alpaca 0.51 0.50 0.52 0.63 0.50 0.50 0.50 0.51 0.51 0.49 0.65 0.74 0.55
Aya 0.57 0.54 0.64 0.67 0.53 0.56 0.57 0.62 0.56 0.61 0.64 0.78 0.61
Bactrian 0.52 0.50 0.53 0.60 0.49 0.51 0.50 0.51 0.51 0.52 0.52 0.71 0.54
sPhinX-T 0.57 0.50 0.64 0.72 0.50 0.50 0.58 0.57 0.57 0.62 0.71 0.83 0.61
sPhinX-0s 0.54 0.5 0.58 0.63 0.51 0.55 0.52 0.52 0.54 0.57 0.64 0.8 0.58
sPhinX missing0.64 0.54 0.73 0.80 0.53 missing0.61 0.59 0.63 0.67 0.66 0.80 0.91 0.68
Phi-3-small
Base Model 0.55 0.51 0.80 0.93 0.52 0.54 0.46 0.56 0.61 0.66 0.86 0.98 0.64
IFT 0.55 0.57 0.81 0.93 0.53 0.58 0.48 0.60 0.62 0.69 0.88 0.96 0.68
m-Alpaca 0.53 0.54 0.80 0.92 0.49 0.54 0.51 0.59 0.64 0.68 0.87 0.99 0.68
Aya 0.60 0.55 0.72 0.83 0.52 0.55 0.52 0.62 0.59 0.69 0.75 0.89 0.65
Bactrian 0.62 0.56 0.83 0.91 0.52 0.60 0.52 0.66 0.65 0.71 0.86 0.98 0.70
sPhinX-T 0.55 0.58 0.83 0.93 0.51 0.57 0.56 0.66 0.68 0.68 0.87 0.98 0.70
sPhinX-0s 0.59 0.58 0.84 0.93 0.50 0.60 0.54 0.63 0.68 missing0.72 0.89 0.96 0.71
sPhinX 0.59 missing0.60 missing0.85 missing0.94 missing0.52 0.57 missing0.58 missing0.68 missing0.69 0.71 missing0.90 missing0.99 missing0.72
Table 10: Granular results for XCOPA (4-shot) on our model. Metric: Accuracy. The best performing IFT dataset for each model is indicated in bold, and the overall best performing IFT model is indicated with an underline.
Language ar en es eu hi id my ru sw te zh avg
Mistral-7B
Base Model 0.65 0.89 0.83 0.56 0.62 0.76 0.52 0.81 0.56 0.52 0.80 0.68
IFT 0.70 0.95 0.92 0.54 0.69 0.79 0.57 0.90 0.58 0.54 0.88 0.73
m-Alpaca 0.53 0.73 0.70 0.51 0.51 0.57 0.50 0.66 0.52 0.52 0.71 0.59
Aya 0.64 0.86 0.81 0.56 0.71 0.73 0.60 0.82 0.67 0.60 0.81 0.71
Bactrian 0.69 0.82 0.74 0.52 0.59 0.76 0.54 0.73 0.62 0.61 0.76 0.67
sPhinX-T 0.78 0.92 0.87 0.56 0.80 0.81 0.66 0.86 0.74 0.70 0.86 0.78
sPhinX-0s 0.57 0.66 0.64 0.47 0.56 0.61 0.50 0.62 0.56 0.52 0.69 0.58
sPhinX 0.83 0.96 0.94 0.57 missing0.84 0.87 missing0.67 0.91 missing0.80 missing0.69 0.94 0.81
Phi-3-small
Base Model 0.80 0.98 0.96 0.61 0.72 0.92 0.53 0.96 0.61 0.55 0.94 0.78
IFT 0.81 0.98 0.96 0.61 0.75 0.92 0.56 0.96 0.61 0.53 0.94 0.79
m-Alpaca 0.81 0.98 0.98 0.58 0.76 0.93 0.52 0.97 0.64 0.54 0.96 0.79
Aya 0.77 0.98 0.97 0.57 0.77 0.93 0.53 0.96 0.74 0.56 0.94 0.79
Bactrian 0.83 0.98 0.98 0.61 0.83 0.94 0.54 0.97 0.79 0.63 0.94 0.82
sPhinX-T 0.84 0.98 0.98 0.60 0.80 0.95 0.52 0.96 0.72 0.60 0.85 0.80
sPhinX-0s 0.84 0.98 0.97 missing0.64 0.77 0.95 0.52 0.96 0.74 0.57 0.95 0.81
sPhinX missing0.86 missing0.99 missing0.99 0.61 0.82 missing0.96 0.54 missing0.98 0.74 0.61 missing0.97 missing0.82
Table 11: Granular results for XStoryCloze (4-shot) on our model. Metric: Accuracy. The best performing IFT dataset for each model is indicated in bold, and the overall best performing IFT model is indicated with an underline.
Language ar de es en fi fr hi it jp ko ta te vi zh avg
Mistral-7B
Base Model 0.25 0.23 0.23 0.24 0.23 0.23 0.26 0.24 0.26 0.23 0.23 0.25 0.26 0.25 0.24
IFT 0.32 0.60 0.62 0.74 0.36 0.62 0.32 0.61 0.43 0.47 0.27 0.27 0.39 0.58 0.47
m-Alpaca 0.32 0.50 0.53 0.56 0.45 0.51 0.27 0.51 0.40 0.41 0.26 0.26 0.33 0.48 0.41
Aya 0.34 0.43 0.43 0.48 0.38 0.47 0.35 0.44 0.4 0.36 0.27 0.25 0.37 0.42 0.38
Bactrian 0.24 0.27 0.25 0.25 0.26 0.27 0.24 0.28 0.26 0.26 0.23 0.23 0.34 0.28 0.26
sPhinX-T 0.66 0.74 0.71 0.81 0.66 0.76 0.57 0.73 0.66 0.68 0.54 0.46 0.66 0.71 0.68
sPhinX-0s 0.64 0.75 0.75 0.82 0.66 0.79 0.53 0.73 0.69 0.66 0.48 0.44 0.66 0.75 0.67
sPhinX 0.69 0.80 0.69 0.87 0.71 0.82 missing0.60 0.79 0.73 0.73 missing0.56 missing0.48 0.70 0.80 0.71
Phi-3-small
Base Model 0.54 0.87 0.85 0.92 0.58 0.86 0.41 0.86 0.70 0.58 0.26 0.30 0.62 0.82 0.65
IFT 0.63 0.89 0.88 0.93 0.63 0.88 0.48 0.88 0.77 0.68 0.32 0.32 0.68 0.85 0.70
m-Alpaca 0.65 0.92 0.90 0.94 0.74 0.91 0.54 0.90 0.80 0.70 0.47 0.45 0.72 0.84 0.75
Aya 0.58 0.86 0.85 0.91 0.65 0.87 0.50 0.86 0.76 0.67 0.37 0.35 0.69 0.84 0.70
Bactrian 0.67 0.88 0.88 0.92 0.70 0.88 0.51 0.86 0.77 0.70 0.37 0.37 0.74 0.86 0.72
sPhinX-T 0.71 0.90 0.90 0.93 0.73 0.90 0.56 0.89 0.82 0.75 0.42 0.38 0.75 0.87 0.75
sPhinX-0s 0.73 0.91 0.90 0.93 0.75 0.92 0.57 0.91 0.82 missing0.82 0.45 0.40 0.76 0.89 0.77
sPhinX missing0.74 missing0.93 missing0.91 missing0.94 missing0.77 missing0.93 0.58 missing0.92 missing0.84 0.76 missing0.46 0.40 missing0.78 missing0.89 missing0.79
Table 12: Granular results for Belebele (0-shot) on our model. Metric: Accuracy. The best performing IFT dataset for each model is indicated in bold, and the overall best performing IFT model is indicated with an underline.
Language ar fr de ro ja ru zh avg
Mistral-7B
Base Model 0.48 0.63 0.65 0.56 0.43 0.54 0.48 0.54
IFT 0.37 0.61 0.61 0.58 0.28 0.5 0.49 0.49
m-Alpaca 0.34 0.51 0.51 0.46 0.29 0.36 0.41 0.41
Aya 0.30 0.50 0.51 0.46 0.24 0.35 0.41 0.39
Bactrian 0.40 0.55 0.52 0.51 0.34 0.44 0.42 0.45
sPhinX-T 0.39 0.60 0.60 0.54 0.39 0.49 0.48 0.49
sPhinX-0s 0.40 0.57 0.61 0.53 0.40 0.51 0.44 0.49
sPhinX 0.45 0.63 0.64 0.59 0.45 0.55 0.52 0.54
Phi-3-small
Base Model 0.48 0.61 0.65 0.57 0.45 0.52 0.53 0.54
IFT 0.43 0.62 0.63 0.52 0.41 0.49 0.50 0.54
m-Alpaca 0.44 0.60 0.61 0.56 0.32 0.44 0.19 0.45
Aya 0.44 0.56 0.57 0.54 0.12 0.45 0.17 0.41
Bactrian 0.48 0.63 0.65 0.59 0.45 0.52 0.52 0.54
sPhinX-T 0.48 0.64 0.65 0.59 0.46 0.53 0.52 0.55
sPhinX-0s 0.49 0.63 0.65 0.59 0.46 0.54 0.53 0.56
sPhinX 0.49 0.64 0.66 0.60 0.46 0.54 0.53 0.56
Table 13: Granular results for Translation for language to English direction (4-shot) on our model. Metric: ChrF. The best performing dataset for each model is indicated in bold, and the overall best performing model is indicated with an underline.
Language ar fr de ro ja ru zh avg
Mistral-7B
Base Model 0.29 0.60 0.54 0.52 0.21 0.34 0.47 0.42
IFT 0.16 0.58 0.57 0.48 0.17 0.29 0.43 0.38
m-Alpaca 0.15 0.62 0.66 0.40 0.12 0.48 0.44 0.41
Aya 0.12 0.54 0.63 0.48 0.16 0.39 0.42 0.39
Bactrian 0.14 0.51 0.49 0.44 0.10 0.35 0.40 0.38
sPhinX-T 0.31 0.58 0.53 0.47 0.20 0.30 0.26 0.38
sPhinX-0s 0.30 0.60 0.55 0.51 0.22 0.31 0.45 0.42
sPhinX 0.35 0.61 0.60 0.55 0.26 0.36 0.49 0.46
Phi-3-small
Base Model 0.31 0.63 0.60 0.43 0.24 0.30 0.45 0.42
IFT 0.29 0.61 0.58 0.39 0.21 0.27 0.41 0.46
m-Alpaca 0.39 0.60 0.58 0.31 0.11 0.24 0.47 0.38
Aya 0.17 0.62 0.56 0.45 0.22 0.28 0.46 0.39
Bactrian 0.30 0.60 0.56 0.47 0.18 0.24 0.44 0.40
sPhinX-T 0.33 0.63 0.60 0.48 0.24 0.31 0.48 0.43
sPhinX-0s 0.32 0.63 0.61 0.50 0.27 0.36 0.49 0.45
sPhinX 0.35 0.64 0.61 0.51 0.26 0.33 0.49 0.46
Table 14: Granular results for Translation for English to language direction (4-shot) on our model. Metric: ChrF. The best performing dataset for each model is indicated in bold, and the overall best performing model is indicated with an underline.
Code Languages Script Data
af Afrikaan Latin 20206
ar Arabic Arabic 26803
bn Bengali Bengal 20165
bg Bulgarian Cyrillic 17300
my Burmese Burmese 12123
zh-Hans Chinese_Simplified Han 100650
zh-Hant Chinese_Traditional Hant 32363
hr Croatian Latin 17340
cs Czech Latin 32711
da Danish Latin 36348
nl Dutch Latin 36586
en English Latin 199900
et Estonian Latin 17207
fi Finnish Latin 33622
fr French Latin 100337
de German Latin 100265
el Greek Greek 17317
he Hebrew Hebrew 24483
hi Hindi Devanagari 20240
hu Hungarian Latin 31999
is Icelandic Latin 20164
id Indonesian Latin 17297
it Italian Latin 85175
jp Japanese Japanese 98366
ko Korean Hangul 30890
lv Latvian Latin 17247
lt Lithuanian Latin 17232
ml Malayalam Malayalam 19817
mr Marathi Devanagari 20069
ne Nepali Devanagari 20092
nb Norwegian Latin 36811
or Oriya Oriya 19153
pl Polish Latin 34711
pt Portuguese Latin 37229
pa Punjabi Gurmukhi 20026
ro Romanian Latin 17149
ru Russian Cyrillic 20108
sr Serbian Latin 17165
sk Slovak Latin 17255
sl Slovenian Latin 17300
es Spanish Latin 100351
sw Swahili Latin 20170
sv Swedish Latin 36533
ta Tamil Tamil 19807
te Telugu Telugu 19947
th Thai Thai 17322
tr Turkish Latin 34405
uk Ukrainian Cyrillic 17282
ur Urdu Perso-Arabic 20162
vi Vietnamese Latin 17358
cy Welsh Latin 20207
Table 15: Language Distribution in Sphinx Dataset